# Spatio-Temporal Foundation Models: Vision, Challenges, and Opportunities

### v1

@@ Abstract
Foundation models have revolutionized artificial intelligence, setting new benchmarks in performance and enabling transformative capabilities across a wide range of vision and language tasks. However, despite the prevalence of spatio-temporal data in critical domains such as transportation, public health, and environmental monitoring, spatio-temporal foundation models (STFMs) have not yet achieved comparable success. In this paper, we articulate a vision for the future of STFMs, outlining their essential characteristics and the generalization capabilities necessary for broad applicability. We critically assess the current state of research, identifying gaps relative to these ideal traits, and highlight key challenges that impede their progress. Finally, we explore potential opportunities and directions to advance research towards the aim of effective and broadly applicable STFMs.
@@ I Introduction
Figure 1: Different types of spatio-temporal data and a selection of related applications.
The emergence of deep learning has significantly advanced state-of-the-art performance in an astonishingly diverse range of applications. In recent years, foundation models (FMs) [8], large neural networks pre-trained on large-scale and broad data, have enabled transformative success particularly in language and vision tasks due to their remarkable abilities to “generalize” to a wide range of downstream tasks through the notion of transfer learning. However, FMs have yet to achieve a similar impact on state-of-the-art performance in tasks involving spatio-temporal data. Encompassing all kinds of data with both spatial and temporal dimensions, spatio-temporal (ST) data is pervasive across an extremely diverse range of fields, including urban analysis [53, 37, 50, 58, 45], weather forecasting [38, 9, 30], climate science [54, 18, 31, 16], environmental monitoring [26, 1, 4, 51], agriculture [34, 10, 14, 55], and public health [42, 36, 60, 49, 64]. As the collection and distribution of spatio-temporal data continues to grow from diverse sources, so does the feasibility and potential of spatio-temporal foundation models (STFMs) to learn shared patterns across different domains, improving efficiency and enhancing generalization, particularly for data-deficient applications. However, progress remains slow due to a number of characteristics of ST data which significantly complicates the learning of STFMs. Moreover, current research is highly fragmented by application, which hinders progress towards a truly universal STFM comparable to those developed for language and vision.
In this paper, we present a vision for the future of STFMs and their potential to advance state-of-the-art performance across diverse spatio-temporal applications. We outline the key generalization capabilities that are essential for broadly applicable STFMs and examine the primary challenges and obstacles to their development. We critically evaluate the current state of research in this direction, identifying gaps relative to these ideal traits. We also explore opportunities for further advancement through targeted research and innovation. In summary, the key contributions of this work are as follows:
@@ II Preliminaries
II-A Spatio-Temporal Data
Spatio-temporal data is any kind of data that involves both spatial and temporal dimensions. Most abstractly, it can be understood as a set of time-series sequences, each of which is associated with a specific spatial location. We denote a ST sequence as $X\in\mathbb{R}^{V\times N\times T}$, where $V$ represents the number of variables or features, $N$ is the number of spatial locations (typically but not necessarily determined by the number of sensors or measuring devices), and $T$ is the number of time-steps. In practice, ST data comes in several different formats or structures, and different formats are suited to different applications. Figure 1 categorizes the four types of ST data, along with illustrative examples of their applications. We now describe each type of ST data in detail.
@@ II-A Spatio-Temporal Data
Raster data is structured on a regular and fixed spatial grid with height $H$ and width $W$. In this format, $N$ is equal to the total number of cells in the grid, i.e. $N=H\times W$. Note that the features in each cell may not correspond to a unique data recording or sensor within the physical system, which depends on physical and logistical constraints. Instead, raw measurements can be converted into raster data of a desired resolution through various interpolation techniques. Raster data is widely used in ST applications, ranging from transportation, weather & climate analysis, medical imaging, remote sensing, and many more. Video data can also be considered a special form of raster data, in which each cell represents a pixel and contains its own set of RGB values at each time-step.
Point reference data is similar to raster data except the spatial locations of data measurements are subject to change from one time step to another. For example, climate data collected by weather balloons which move over time due to air currents, or sea surface temperatures recorded by sensors attached to drifting buoys. In this case, the evolving spatial position becomes an additional variable that must be tracked and recorded at each time step, adding complexity to the data structure.
@@
Trajectories represent the paths traced by moving objects through space over time, consisting of paired sets of geographical co-ordinates and time-stamps: $\{l_{i},t_{i}\}$. They are commonly found in applications related to mobility, such as pedestrian or vehicular movement tracking. When dealing with multiple moving objects, it is typical to group trajectories into discrete buckets, where the features represent the number of trajectories within a specific spatial boundary over a specific time-period, closely resembling raster data.
@@
Event Data is characterized by a set of tuples $\{e_{i},l_{i},t_{i}\}$, where each tuple corresponds to a discrete event of type $e_{i}$, recorded at location $l_{i}$ and time $t_{i}$. Events are typically rare, such as crimes or traffic accidents. As a result, event data is typically much sparser than other forms of ST data, with a large proportion of zero entries. This sparsity presents unique challenges and requires specialized techniques to model and analyze the data effectively.
@@
As outlined in [20], spatio-temporal data exhibits two key properties. The first property is heterogeneity, which means that ST patterns can vary across spatial (from one location to another) and temporal (from one time period to another) ranges and scales. Heterogeneity is a particularly challenging property, as it violates the assumption that all data samples are independent and identically distributed, i.e. drawn from the same probability distribution. The second property is auto-correlation, which reflects that measurements taken closer together tend to follow more similar distributions, where closeness can be understood in both the spatial and the temporal sense. This is neatly summarized in Tobler’s First Law of Geography, which states that “everything is related to everything else, but near things are more related than distant things”.
II-B Spatio-Temporal Data Mining
@@
Figure 2: Illustration of the four types of generalization. Top-left: domain generalization - across different sources of data representing different physical systems and categories of applications. Top-right: spatial generalization - across different locations or regions in space. Bottom-left: temporal generalization - across different periods and intervals of time. Bottom-right: scale generalization - across different resolutions, frequencies or granularities of data.
@@
Spatio-temporal data mining involves learning to model both the spatial and temporal patterns within a set of ST data. In recent years, neural networks that integrate convolutional and recurrent modules have proven particularly successful at capturing both spatial and temporal dependencies. CNN-LSTMs firstly use a convolutional neural network (CNN) to extract spatial features from input data, followed by a long short-term memory (LSTM) network to learn sequential patterns from the extracted spatial features [7, 57, 11]. Alternatively, ConvLSTMs [41, 27, 5, 22, 47, 15] replace the matrix multiplications in the LSTM gates with convolution operations to capture spatial dependencies within the sequential model.
@@
Spatio-temporal graph neural networks (ST-GNNs) [39, 21, 44, 33, 58, 3, 30, 65, 40] have recently gained prominence due to their ability to flexibly handle ST data that does not conform to a regular grid structure. Instead, they operate over spatio-temporal graphs, which represent the spatial locations as nodes or vertices in a graph, with connecting edges representing spatial relationships between adjacent nodes, such as proximity or connectivity. More details about ST-GNNs can be found in [20].
Figure 3: A selection of spatio-temporal applications that an STFM should be able to generalize across. Dashed lines indicate the potential for correlations or relationships in the exhibited patterns of different applications.
@@
Following success in other modalities, transformers [46] have also gained significant attention for spatio-temporal data [56, 2, 59, 23, 17, 29] due to their ability to capture long-range dependencies across space and time through the self-attention mechanism. Unlike convolutional models, which operate over local receptive fields, or recurrent models, which depend on sequential processing, transformers can learn global relationships by attending to all parts of an input sequence simultaneously. This is particularly useful in domains where complex, non-linear interactions evolve over large spatial areas and extended time periods. Transformer models are also heavily relied upon in foundation models in other modalities, However, the vast majority of ST research continues to adopt a one-model-per-task paradigm, in which separate models are trained for single tasks with specific training data.
@@ III Spatio-Temporal Foundation Models
  • In-distribution generalization: We can assess the model’s ability to learn generalizable patterns across multiple domains and distributions during pre-training by evaluating its performance on tasks encountered during that phase.
  • Out-of-distribution generalization: Alternatively, we can evaluate the model’s ability to transfer to new tasks that were not observed during pre-training, even without prior exposure to their data distribution.
@@ III-A1 Domain Generalization
As discussed, ST data is prevalent in a highly diverse range of applications. Figure 3 illustrates a small selection of example applications, organized by top-level categories or domains: transportation, weather & climate and urban activity. Each of these categories encompasses a plethora of individual applications. For example, the transportation domain includes traffic flow measured by road network sensors, the time and location of traffic incidents, or the volume of passengers entering and exiting a public transport network at different stops. Within weather & climate, there is a wide variety of different atmospheric variables, such as air temperature, precipitation, or the concentration of different pollutants. Given these diverse applications, the first type of generalization is across different domains of data.
@@
Foundation models rely on the existence of shared patterns across different data sources or domains. For example, in language, the semantic meaning of words are generally consistent across various contexts, and sentences adhere to a common set of grammatical rules. In contrast, the rules that determine the distribution of ST data are highly application-dependent. For example, in a transportation network, applications like traffic flow and traffic incidents are likely to exhibit strong correlations because they both pertain to the same underlying physical system. Similarly, within weather & climate studies, the concentration of pollutants in the air is often closely correlated with rainfall occurrence and intensity. In these cases, training a spatio-temporal foundation model (STFM) with data from both applications could offer mutual benefits by leveraging these shared patterns. However, the existence of shared patterns becomes much less clear between more disparate applications, such as traffic incidents and disease outbreaks. In such cases, it is uncertain whether training an STFM to simultaneously model both applications would enhance or degrade performance for either, a phenomenon known as negative transfer in the deep learning community. This highlights the challenges of developing an STFM that can generalize effectively across diverse applications. As discussed in Section IV, current research typically focuses on more narrowly defined STFMs, often limited to a few applications within a single domain, rather than addressing cross-domain generalization.
@@ III-A2 Spatial Generalization
The second type of generalization is across different locations in space. An STFM should not only be limited to applications from a limited selection of geographic spaces; it should learn from diverse environments and conditions simultaneously, and also be able to transfer to unseen places during inference.
@@
ST data can exhibit significant spatial heterogeneity. In other words, data patterns can vary significantly from one location to another, even within a single application. For example, in traffic flow, a model trained on traffic data from one city may struggle to generalize to another city with different road networks or traffic regulations. In pollutant concentration, a model trained on data from heavily urbanized areas may struggle to generalize to suburban or rural areas. This challenge is particularly acute in applications where existing dataset available for use in pre-training are heavily biased towards certain regions. For example, a disproportionate amount of the traffic datasets commonly used in existing studies are collected in a small number of major urban centers like Beijing, New York City, and London. This limitation increases the risk that an STFM is biased towards the patterns found in these cities and others like them, while failing to generalize to the vast majority of other regions with a smaller or no presence at all in the training data.

### v2

@@ Abstract
Foundation models have revolutionized artificial intelligence, setting new benchmarks in performance and enabling transformative capabilities across a wide range of vision and language tasks. However, despite the prevalence of spatio-temporal data in critical domains such as transportation, public health, and environmental monitoring, spatio-temporal foundation models (STFMs) have not yet achieved comparable success. In this paper, we articulate a vision for the future of STFMs, outlining their essential characteristics and the generalization capabilities necessary for broad applicability. We critically assess the current state of research, identifying gaps relative to these ideal traits, and highlight key challenges that impede their progress. Finally, we explore potential opportunities and directions to advance research towards effective and broadly applicable STFMs.
@@ I Introduction
Figure 1: The four types of spatio-temporal data (raster, point reference, trajectory and events), with example use cases and illustrations.
The emergence of deep learning has significantly advanced state-of-the-art performance in an astonishingly diverse range of applications. In recent years, foundation models (FMs) [8], which are large neural networks pre-trained on large-scale and broad data, have enabled transformative success due to their powerful abilities to generalize to a wide range of downstream tasks through the notion of transfer learning. Models like GPT [9], Llama [60], and BERT [18] and DeepSeek [22] are trained on massive natural language corpora, enabling them to capture nuanced linguistic patterns which are important for various downstream tasks such as text classification, machine translation, summarization, and question answering. Similarly, models for computer vision such as CLIP [50], ALIGN [25] and SAM [31] have greatly improved performance across a wide variety of vision domains and tasks, such as classification, object detection and segmentation. There has also been very recent progress in foundation models for time-series data [35] with models like Lag-llama [51], Time-LLM [27] and Moirai [67], where time-series tasks vary between domains and temporal frequencies. However, despite this success, FMs have yet to achieve a similar impact on performance in spatio-temporal tasks. Spatio-temporal (ST) data, encompassing data with both spatial and temporal dimensions, is pervasive across an extremely diverse range of fields, including urban analysis [69, 49, 65, 74, 59], weather forecasting [52, 10, 41], climate science [70, 23, 43, 20], environmental monitoring [36, 1, 4, 66], agriculture [46, 12, 17, 71], and public health [56, 48, 76, 64, 80]. Unlike time-series, ST data exhibit both spatial and temporal dependencies, which both need to be captured simultaneously. As the collection and distribution of spatio-temporal data continues to grow from diverse sources, so does the feasibility of spatio-temporal foundation models (STFMs) to learn shared patterns across different domains. This has great potential to improving efficiency, enhance generalization, particularly for data-deficient applications, and improve performance in a wide range of related tasks. However, progress remains slow due to the characteristics of ST data which makes training STFMs particularly challenging. Moreover, we observe that current research in this direction is highly fragmented between applications, which hinders progress towards a truly universal STFM comparable to those developed for other modalities.
In this paper, we present a vision for STFMs and aim to unify spatio-temporal foundation model research. We outline the key generalization capabilities that are essential for broadly applicable STFMs and examine the primary challenges and obstacles to their development. We critically evaluate the current state of research in this direction, identifying gaps relative to these ideals, and we explore opportunities for further advancement through targeted research and innovation. In summary, the key contributions of this work are as follows:
@@ II Preliminaries
II-A Spatio-Temporal Data Types
Spatio-temporal data is any kind of data that involves both spatial and temporal dimensions. In practice, ST data can be organized into one of four main types or categories: raster, point-reference, trajectory and event data. Figure 1 shows illustrations and example applications of each type of ST data, and we now describe them in more detail.
@@ II-A Spatio-Temporal Data Types
Raster data is ST data that is organized into a regular and fixed grid of cells, where each cell corresponds to a specific spatial location and contains a time-evolving sequence of measurements. With this, the data structure is represented as a four dimensional matrix: $X\in\mathbb{R}^{C\times L\times H\times W}$, where $C$ is the number of measured variables (or features), $L$ is the length of the sequence (i.e. number of time-steps), $H$ the height and $W$ the width of the grid. The total number of cells is $H\times W$. In practice, each cell may not correspond to a unique recording or sensor, as these are determined by physical or logistical constraints in the measured system. In this case, raw measurements are often processed into raster data of a desired resolution through various pooling and interpolation techniques. Due to its uniform structure and relative simplicity, raster data is widely used in various ST applications, ranging from transportation, weather & climate analysis, medical imaging, remote sensing, and more. Video can be seen as a special case of raster data, where the measured variables are the RGB values in each pixel and each time-step corresponds to a frame. In this paper, we omit discussion of video data and their related applications, as their information-rich nature and complex pixel-level visual dynamics diverge significantly from the majority of spatio-temporal applications and are more appropriately aligned with computer vision.
Point reference data are ST sequences measured at specific point co-ordinates and may not neatly form a uniform grid. In this case, the data is represented as a three dimensional matrix: $X\in\mathbb{R}^{C\times L\times N}$, where $C$ represents the number of features, $L$ the sequence length and $N$ the number of points, which is typically (but not necessarily) equal to the number of sensors or measuring devices. These positions may be fixed, for example weather stations which are fixed but not uniformly distributed through space. On the other hand, other devices such as weather balloons or sensors attached to ocean buoys are subject to atmospheric and ocean currents, meaning their position is subject the change from one time-stamp to another.
Figure 2: A framework for spatio-temporal foundation models (STFMs). Top-left: STFMs can flexibly handle various forms of ST data as input (see Section II). Middle-left: Novel techniques for training STFMs to handle multiple domains. Bottom-left: Various forms of complementary information in a variety of modalities can be incorporated into STFMs as guidance to perform specific tasks. Top-right: the generalization capabilities expected of STFMs (see Section III-A. Middle-right: A variety of relevant applications for STFMs. Bottom-right: Diverse types of tasks that STFMs should perform.
@@
Trajectories represent the paths traced by moving objects through space over time, consisting of paired sets of geographical co-ordinates and time-stamps: $\{l_{i},t_{i}\}_{i=1:L}$. They are more suited to applications related to mobility, such as human, animal, vehicle or tropical cyclone tracking. In applications that involve several objects, it is common to discretize trajectories into buckets that represent a fixed spatial and temporal boundary, and the measured features indicate the number of trajectories found in each bucket. In this case, trajectory data is processed to more closely resemble raster data.
@@
Event Data is characterized by a set of tuples $\{e_{i},l_{i},t_{i}\}_{i=1:N}$, where each tuple corresponds to a discrete event of type $e_{i}$, recorded at location $l_{i}$ and time $t_{i}$. Common applications include occurrences such as crime incidents, traffic accidents, disease outbreaks, natural disasters, and more. Event data can be particularly sparse, with large proportion of zero entries (i.e. no events within a given spatial and temporal interval), which presents unique challenges that require specialized techniques.
@@
[26] outlines two key properties shared by all kinds of spatio-temporal data, these are heterogeneity and auto-correlation. Heterogeneity means that ST patterns can vary across spatial (from one location to another) and temporal (from one time period to another) ranges and scales. This is a particularly challenging property for deep learning, as it is a violation of the i.i.d. assumption (that all data are independent drawn from an identical distribution), which is fundamental to the effective training of neural networks. Secondly, auto-correlation means that measurements taken closer together tend to follow more similar distributions, where closeness is understood in both the spatial and the temporal sense. This is neatly summarized in Tobler’s First Law of Geography, which states that “everything is related to everything else, but near things are more related than distant things”.
II-B Spatio-Temporal Tasks
Prediction is the task of forecasting future states or events based on historical patterns observed in spatio-temporal data. This process typically involves two components: the window, which represents the set of historical observations used as input, and the horizon, which defines the future time period to be predicted. In some cases, there can be a delay between the window and the horizon intervals, however it is more common to assume temporal continuity from the window to the horizon to maintain temporal continuity. Also, it is typical to assume that the window and the horizon contain an equal number of time steps. Prediction tasks can be categorized into short-term and long-term prediction, depending on the number of time steps involved, with short-term predictions focusing on immediate future states and long-term predictions concerning more distant outcomes, which are typically more challenging. Likely due to its practical importance in a variety of domains, spatio-temporal prediction has attracted the greatest amount of interest from the research community.
Classification tasks involve assigning labels to ST sequences out of a set of candidate labels. Examples include identifying land use from satellite imagery, classifying weather events, or vehicle types based on their trajectories.
Clustering aims to group spatio-temporal data points based on similarity. This is an important task for identifying and extracting relationships between different entities. Here, similarity can be flexibly defined to suit the application in question e.g. segmenting regions with similar temperature trends or identifying hotspots of disease outbreaks.
Anomaly Detection involves identifying spatio-temporal patterns that deviate significantly from the normal or expected patterns, such as detecting abnormal rainfall patterns or unusual movements in urban mobility. Anomalies are generally assumed to be rare in comparison to normal data, therefore anomaly detection methods often build a model of the expected behavior of a certain instance in order to detect any discrepancy to the norm.
Change Detection is the process of detecting significant deviations in spatio-temporal data over time. This is similar to anomaly detection, however anomalies are generally assumed to be isolated incidents, whereas change detection is aimed towards detecting sustained changes to the underlying environment, such as monitoring deforestation using satellite images or analyzing urban expansion.
Imputation tasks focus on filling missing values in spatio-temporal datasets, such as gaps in weather station readings or satellite observations. Missing values can be temporal in nature, i.e. a missing time-step from an otherwise complete sequence, or spatial in nature, i.e. an entire sequence for an unobserved spatial location.
II-C Spatio-Temporal Data Mining
@@
Figure 3: Four forms of generalization in spatio-temporal data. Top-left: domain generalization across different sources of data representing different physical systems and categories of applications. Top-right: spatial generalization: across different locations or regions in space. Bottom-left: temporal generalization across different periods and intervals of time. Bottom-right: scale generalization across different resolutions, frequencies or granularities of data.
@@
Spatio-temporal data mining involves learning to model both the spatial and temporal patterns within ST data. A common approach has been to develop neural networks that integrate convolutional modules to capture spatial dependencies with recurrent modules to capture temporal dependencies. CNN-LSTMs use a convolutional neural network (CNN) to extract spatial features from input data, followed by a long short-term memory (LSTM) network to learn sequential patterns from the extracted spatial features [7, 73, 13]. Alternatively, ConvLSTMs [55, 37, 5, 30, 62, 19] replace the matrix multiplications in the LSTM gates with convolution operations to capture spatial dependencies within the sequential model.
@@
In recent years, spatio-temporal graph neural networks (ST-GNNs) [53, 28, 58, 45, 74, 3, 41, 81, 54] have gained prominence due to their ability to flexibly handle ST data that does not conform to a regular grid structure. Instead, they operate over spatio-temporal graphs, which represent spatial positions as nodes or vertices in a graph, with connecting edges representing spatial relationships between adjacent nodes, such as proximity or connectivity. More details about ST-GNNs can be found in [26].
@@
Following success in other modalities, the transformer architecture [61] has seen significant success in spatio-temporal applications [72, 2, 75, 32, 21, 40] due to their ability to capture long-range dependencies across space and time through the self-attention mechanism. Unlike convolutional models, which operate over local receptive fields, or recurrent models, which depend on sequential processing, transformers can learn global relationships by attending to all parts of an input sequence simultaneously. This is particularly useful in domains where complex, non-linear interactions evolve over large spatial areas and over long temporal periods. Transformers are also the architecture of choice for foundation models for both image and language modalities, though the vast majority of ST research continues to adopt a one-model-per-task paradigm which requires training separate models for individual tasks.
@@ III Spatio-Temporal Foundation Models
Figure 4: A selection of spatio-temporal domains that STFMs should be able to generalize across. Dashed lines indicate examples of potential for correlations or relationships in data patterns between different domains.
@@
  • In-distribution (ID) generalization: We can assess the model’s ability to learn across multiple domains and distributions simultaneously during the pre-training stage by evaluating its performance on the variety of tasks seen during pre-training.
  • Out-of-distribution (OOD) generalization: Alternatively, we can evaluate the model’s ability to transfer to new tasks that were not observed during pre-training, even without prior exposure to their data distribution.
@@ III-A1 Domain Generalization
As we have seen, ST data is prevalent in a highly diverse range of applications. Figure 4 illustrates a small selection of these, organized by top-level categories or domains: transportation, weather & climate and urban activity. Each of these categories encompasses a plethora of individual applications. For example, the transportation domain includes traffic flow measured by road network sensors, the time and location of traffic incidents, or the volume of passengers entering and exiting a public transport network at different stops. Within weather & climate, there is a wide variety of different atmospheric variables, such as air temperature, precipitation, or the concentration of different pollutants. Given these diverse applications, the first type of generalization is across different domains of data.
@@
Foundation models rely on the existence of shared patterns across different data sources or domains. For example, in language, the semantic meaning of words are generally consistent across various contexts, and sentences adhere to a common set of grammatical rules. In contrast, the rules that determine the distribution of ST data are highly application-dependent. For example, in a transportation network, applications like traffic flow and traffic incidents are likely to exhibit strong correlations because they both pertain to the same underlying physical system. Similarly, within weather & climate studies, the concentration of pollutants in the air is often closely correlated with rainfall occurrence and intensity. In these cases, training a spatio-temporal foundation model (STFM) with data from both applications could offer mutual benefits by leveraging these shared patterns. However, the existence of shared patterns becomes much less clear between more disparate applications, such as traffic incidents and disease outbreaks. In such cases, it is uncertain whether training STFMs to simultaneously model both applications would enhance or degrade performance for either, a phenomenon known as negative transfer in the deep learning community. This highlights the challenges of developing STFMs that can generalize effectively across diverse applications. As discussed in Section IV, current research typically focuses on more narrowly defined STFMs, often limited to a few applications within a single domain, rather than addressing cross-domain generalization.
@@ III-A2 Spatial Generalization
The second type of generalization is across different locations in space. STFMs should not only be limited to applications from a limited selection of geographic spaces; it should learn from diverse environments and conditions simultaneously, and also be able to transfer to unseen places during inference.
@@
ST data can exhibit significant spatial heterogeneity. In other words, data patterns can vary significantly from one location to another, even within a single application. For example, in traffic flow, a model trained on traffic data from one city may struggle to generalize to another city with different road networks or traffic regulations. In pollutant concentration, a model trained on data from heavily urbanized areas may struggle to generalize to suburban or rural areas. This challenge is particularly acute in applications where existing dataset available for use in pre-training are heavily biased towards certain regions. For example, a disproportionate amount of the traffic datasets commonly used in existing studies are collected in a small number of major urban centers like Beijing, New York City, and London. This limitation increases the risk that STFMs are biased towards certain patterns that are most prominent in these and other similar cities, and failing to generalize to other regions with little or no representation in the training data.